OpenStudy (anonymous):

examine the differentiability of the function f(x)=x^2+2 at x=2

OpenStudy (dumbcow):

limit of f(x) at x=2 exists and function is continuous so f(x) is differentiable at x=2

OpenStudy (anonymous):

it is differentiable at x=2 f\(x)=2x

OpenStudy (munish):

f'(x)=2x f'(2)=4

OpenStudy (anonymous):

it is polynomial and polynomials are always continuous and (so limit exits) so it means this is differentiable at x=2 f'(x)=2x f'(2)=2(2)=4

OpenStudy (anonymous):

but who i know if differentible or not there is some method to prove

OpenStudy (anonymous):

yes there is for example f(x)=|x| is not differentiable at x=0 and the function f(x)=x^(2/3) is not differentiable at x=0 because take its derivative f'(x)=2/3/(x^1/3) if i put x=0 then the function gets undefined if limit exists then most probably derivative exists

OpenStudy (munish):

a function is differentiable if the limit at that point from left hand and right hand is same and exists.

OpenStudy (anonymous):

thx sami

OpenStudy (anonymous):

i think i want to solve by ur way @munish

OpenStudy (anonymous):

a differentiable function is a function whose derivative exists at each point in its domain. if x0 is a point in the domain of a function ƒ, then ƒ is said to be differentiable at x0 if the derivative ƒ′(x0) is defined. This means that the graph of ƒ has a non-vertical tangent line at the point (x0, ƒ(x0)), and therefore cannot have a break, bend, or cusp at this point. and rest is same as @sami-21 told

OpenStudy (anonymous):

okay thank u @annas and @sami-21

OpenStudy (anonymous):

as i said no thanks to friends friends are for help :)